# Exchangeability and randomness for infinite and finite sequences

### v1

@@ 4.2 Smaller observation spaces $\mathbf{Z}$
On Kolmogorov’s and Martin-Löf’s log scale, $N^{(K-1)/2}$ becomes $\frac{K-1}{2}\log N$, and differences of $O(\log N)$ are often ignored in the algorithmic theory of randomness; according to Kolmogorov, ‘‘we should not be afraid of logarithms (as well as $O(1)$ terms that we have anyway)’’. [^2] In Russian, “логарифмов не надо бояться, так же как и констант”; recorded by Shen [23, note 12] and translated by the authors of [23, arXiv version]. In this sense exchangeability and being IID nearly coincide for finite sequences as well. However, for $\left|\mathbf{Z}\right|=\infty$ the difference becomes $\Theta(N)$ on the logarithmic scale, which is the largest possible in Kolmogorov’s and Martin-Löf’s binary setting.
@@
$\operatorname{\mathbb{P}^{\mathrm{R}}}(A)=\frac{N!}{(N/2)!^{2}}2^{-N}\leq 1=\operatorname{\mathbb{P}^{\mathrm{X}}}(A),$
@@ 5 Conclusion
De Finetti’s theorem has many fascinating generalizations and variations, and we can ask similar questions about those. One generalization that is especially close to the subject of this note concerns weighted exchangeability [2], which accounts for a known covariate shift. Many more are provided by the theory of repetitive structures; see, e.g., [18], [27, Part IV], and [3, Chap. 4] (the last book, however, does not use the terminology of repetitive structures).
@@ Acknowledgments
Thanks to the participants of the workshop ‘‘Algorithmic Statistics’’ (Oxford, November 28, 2025) for a useful discussion.

### v2

@@ 4.2 Smaller observation spaces $\mathbf{Z}$
On Kolmogorov’s and Martin-Löf’s log scale, $N^{(K-1)/2}$ becomes $\frac{K-1}{2}\log N$, and differences of $O(\log N)$ are often ignored in the algorithmic theory of randomness; according to Kolmogorov, ‘‘we should not be afraid of logarithms (as well as $O(1)$ terms that we have anyway)’’. [^2] In Russian, “логарифмов не надо бояться, так же как и констант”; recorded by Shen [24, note 12] and translated by the authors of [24, arXiv version]. In this sense exchangeability and being IID nearly coincide for finite sequences as well. However, for $\left|\mathbf{Z}\right|=\infty$ the difference becomes $\Theta(N)$ on the logarithmic scale, which is the largest possible in Kolmogorov’s and Martin-Löf’s binary setting.
@@
$\operatorname{\mathbb{P}^{\mathrm{R}}}(A)=\frac{N!}{(N/2)!^{2}}2^{-N}\leq 1=\operatorname{\mathbb{P}^{\mathrm{X}}}(A),$ (21)
@@
Remark 8.
The case of a fixed $K:=\left|\mathbf{Z}\right|$ with variable $N$ corresponds to the bottom left corner of Fig. 1. According to (20) (and the end of Remark 7), the slope of the graph of $\log_{10}C$ in that corner becomes infinite ‘‘under microscope’’ as $N\to\infty$; namely, we expect the slope to grow as $\frac{1}{2}\log_{10}N$. For $N=10^{3}$ and $N=10^{6}$, as used in Fig. 1, this expression gives the slopes of 1.5 and 3, respectively; more precise values given by (21) are 1.598 and 3.098, respectively.
Figure 2: The graphs of $\log_{10}C$ for $\mathcal{N}=10^{3}$ (left panel) and $\mathcal{N}=10^{6}$ (right panel, using $10^{5}$ as the unit for the labels on both axes), as described in text.
Figure 2 complements Fig. 1 by plotting $\log_{10}C$ as function of $N$ for a fixed $K$ explicitly. It shows two ranges for $N$, $N\leq 10^{3}$ in the left panel and $N\leq 10^{6}$ in the right panel; the two panels look similar, and the description to follow is applicable to either. Let $\mathcal{N}\in\{10^{3},10^{6}\}$ be the upper limit of the range of $N$. The base graph shows $\log_{10}C$ in black as function of $N$ for $K\geq\mathcal{N}$. The remaining 9 graphs consist of two pieces, black and coloured; the graphs and their labels in the legends are shown in the same order, top to bottom. Let me describe, for concreteness, the bottom one labelled as $K/\mathcal{N}=0.1$; this description will also be applicable, mutatis mutandis, to the other 8 graphs. The graph corresponds to $K=0.1\,\mathcal{N}$ and consists of two parts: the values for $N\leq K$ are shown in black and the values for $N>K$ in olive. The behaviour of the graph changes drastically after $N=K$, which is marked by using a different colour: the black part grows exponentially fast, while the olive part grows only polynomially fast (albeit for a polynomial of a high degree, namely $\lceil(K-1)/2\rceil\geq 50$ according to (20)).
Let us see why the graphs in Fig. 2 look the way they do. The black graph is the diagonal of the bounding rectangle in the limit $\mathcal{N}\to\infty$. Suppose $K=\frac{k}{10}\,\mathcal{N}$, $k\in\{1,\dots,9\}$, is one of the $K$ marked in Fig. 2. If $nK<N<(n+1)K$, all $n_{k}$ in a balanced split (17) are either $n$ or $n+1$. Incrementing $N$ by 1 to $N+1$ leads to replacing one of the $n_{k}=n$ by $n_{k}=n+1$. The relative increment in the constant $C$ given by (16) is
$\left(\frac{(N+1)^{N+1}}{(N+1)!}\frac{(n+1)!}{(n+1)^{n+1}}\right)/\left(\frac{N^{N}}{N!}\frac{n!}{n^{n}}\right)\\ =\left(1+\frac{1}{N}\right)^{N}\left(1+\frac{1}{n}\right)^{-n}\approx\mathrm{e}\left(1+\frac{1}{n}\right)^{-n}$ (22)
(the $\approx$’’ is justified by $N>K$ and all the $K$ marked in Fig. 2 being large, at least 100).
For $n=1$, the last expression in (22) gives the slope of $\log_{10}\mathrm{e}-\log_{10}2\approx 0.133$ between $K$ and $2K$. This is the slope of the full coloured lines for $K=0.9\,\mathcal{N}$ (blue) to $K=0.5\,\mathcal{N}$ (purple) and the slope of the first straight segment of the other coloured lines (strictly speaking, ‘‘straight’’ should be understood as ‘‘approximately straight’’ because of the $\approx$’’ in (22)). The slope of the following straight segment of the coloured lines for $K=0.4\,\mathcal{N}$ (brown) to $K=0.1\,\mathcal{N}$ (olive) is $\log_{10}\mathrm{e}-2\log_{10}1.5\approx 0.082$. The slope of the following straight segment for the bottom three coloured lines is $\log_{10}\mathrm{e}-3\log_{10}(4/3)\approx 0.059$, etc. It is clear from (22) that the slope tends to 0 when $n\to\infty$, and we can see that the bottom line of Fig. 2 is close to being horizontal on the right. If any of the coloured lines is continued to the right beyond $\mathcal{N}$, it will consist of segments of exponentially fast growth with decreasing growth rates, which will make the overall growth rate polynomial.
@@ 5 Conclusion
De Finetti’s theorem has many fascinating generalizations and variations, and we can ask similar questions about those. One generalization that is especially close to the subject of this paper concerns weighted exchangeability [2], which accounts for a known covariate shift. Many more are provided by the theory of repetitive structures; see, e.g., [19], [28, Part IV], and [3, Chap. 4] (the last book, however, does not use the terminology of repetitive structures).
@@ Acknowledgments
Thanks to the participants of the workshop ‘‘Algorithmic Statistics’’ (Oxford, November 28, 2025) for a useful discussion and to Ioannis Kontoyiannis for informing me about [12] and other related work.
I acknowledge the use of Microsoft Copilot in exploring proof ideas, which I reviewed carefully. I take full responsibility for this paper’s mathematical statements and their proofs.
Appendix A Derivation of $L(1)=1$
In this appendix we check the statement made in Remark 7, which is equivalent to $L(1)=1$. We can simplify the expression for $L^{\prime}(x)$ in (19) as
$L^{\prime}(x)=\ln(n!)+n(n+1)\ln\frac{n+1}{n}-n\ln(n+1).$
Integrating $L^{\prime}$ from 0 to 1 gives
$L(1)=\sum_{n=1}^{\infty}\left(\frac{\ln(n!)}{n(n+1)}+\ln\frac{n+1}{n}-\frac{\ln(n+1)}{n+1}\right).$
We are required to show that the partial sums
$S_{N}:=\sum_{n=1}^{N}\frac{\ln(n!)}{n(n+1)}+\sum_{n=1}^{N}\ln\frac{n+1}{n}-\sum_{n=1}^{N}\frac{\ln(n+1)}{n+1}$ (23)
converge to 1 as $N\to\infty$. By telescoping, we can transform the first addend in (23) as
$\sum_{n=1}^{N}\frac{\ln(n!)}{n(n+1)}=\sum_{n=1}^{N}\sum_{k=1}^{n}\frac{\ln k}{n(n+1)}=\sum_{k=1}^{N}\sum_{n=k}^{N}\frac{\ln k}{n(n+1)}\\ =\sum_{k=1}^{N}\ln k\left(\frac{1}{k}-\frac{1}{N+1}\right)=\sum_{k=1}^{N}\frac{\ln k}{k}-\frac{1}{N+1}\sum_{k=1}^{N}\ln k$
(this uses $\frac{1}{n(n+1)}=\frac{1}{n}-\frac{1}{n+1}$ in the third equality) and the second addend as
$\sum_{n=1}^{N}\ln\frac{n+1}{n}=\ln(N+1).$
Plugging this into (23) gives
$S_{N}$ $:=\sum_{k=1}^{N}\frac{\ln k}{k}-\frac{1}{N+1}\sum_{k=1}^{N}\ln k+\ln(N+1)-\sum_{n=1}^{N}\frac{\ln(n+1)}{n+1}$
$=-\frac{\ln(N+1)}{N+1}-\frac{\ln(N!)}{N+1}+\ln(N+1)\sim 1,$
where the ‘‘==’’ is obtained by combining the first and last addends in the preceding expression and the definition of $N!$, and the $\sim$’’ is obtained from Stirling’s formula in the crude form $\ln(N!)=N\ln N-N+O(\ln N)$.